on a frictionless surface. a boy pushes a girl, causing a girl to accelerate at 3.2 m/s to the right. mb= 72 kg; mg=63 kg. find F exerted by boy on girl and F exerted by girl on boy. Find the acceleration of the boy
1
Expert's answer
2019-10-14T09:39:51-0400

From the third Newtons' law:


"F_{bg}=F_{gb}=F"

From the second Newtons' law:


"F=m_ga_g=(63)(3.2)=200\ N"

The acceleration of the boy:


"a_b=\frac{F}{m_b}=\frac{(63)(3.2)}{72}=2.8\frac{m}{s^2}"